jabez177 (jabez177):

I have 10 of these, can someone make sure of the ones I have already answered? And help me with the rest.

jabez177 (jabez177):

Here

jabez177 (jabez177):

The other 5

jabez177 (jabez177):

@melons

OpenStudy (anonymous):

#2 is d=-4+2/5c. You needed to subtract 2c from both sides, not add. I'll check the rest as well.

OpenStudy (anonymous):

2,3,4=wrong

OpenStudy (anonymous):

#3 is actually correct. -6p-4q=18 -6p=18+4q p=-3-2/3q

jabez177 (jabez177):

Okay, you can also click on the "Equation" button where you are typing to type in fractions and expressions, etc. :)

OpenStudy (anonymous):

#4 is -6p-4q=18 -4q=6p+18 q=-3/2q-9/2

OpenStudy (anonymous):

\[2X-Y=6\]\[-y=-2x+6\]\[y=2x-6\]

OpenStudy (anonymous):

Try out the rest and see what you get. I'll help you with any mistakes.

jabez177 (jabez177):

Okay. Can you type the answers out for the other ones that I have answered and the correct answers while I answer the other 6?

OpenStudy (anonymous):

http://www.cymath.com try this site :)

jabez177 (jabez177):

#4 q = -6

OpenStudy (anonymous):

So far we have 1. Correct 2. d = 2/5x - 4 3. Correct 4. q=-3/2q - 9/2 5. Correct 6. y=2x-6. Is that what you're asking me to do?

OpenStudy (anonymous):

Sorry, the x in #2 should be c.

jabez177 (jabez177):

yes

jabez177 (jabez177):

Okay, let me do the rest and can you make sure they're correct using the equation button? Thanks! :)

OpenStudy (anonymous):

\[a=-\frac{ 3 }{4}b+2\]\[4a=-3b+3\]\[4a-8=-3b\]\[-\frac{ 4 }{ 3 }a+\frac{ 8 }{ 3 }=b\]

OpenStudy (anonymous):

\[2a-3bc=-9\]\[2a=3bc-9\]\[a=\frac{ 3 }{ 2}bc-\frac{ 9 }{ 2 }\]

jabez177 (jabez177):

Okay, can you number them. I'm working on #6.

OpenStudy (anonymous):

My latest answer was for #8. The answer before that is for #7.

jabez177 (jabez177):

Okay, thanks and #6 is y = 2x - 6

OpenStudy (anonymous):

Correct. #9: \[2a-3bc=-9\]\[-3bc=-2a-9\]\[b=\frac{ -9-2a }{ -3c }\]

OpenStudy (anonymous):

#10: \[2a-3bc=-9\]\[-3bc=-2a-9\]\[c=\frac{ -9-2a }{ -3b }\]

jabez177 (jabez177):

#10 \[c=\frac{ 2a }{ 3b } + \frac{ 3 }{ b }\]

jabez177 (jabez177):

wait what?

OpenStudy (anonymous):

I'm pretty sure your answer is equivalent to mine. I wouldn't worry.

jabez177 (jabez177):

Okay, so an you number all of them like #1 and the answer then #2 and the answer? This way i can verify accurately that my answers are correct because I somewhat got mixed up. I really appreciate it!

OpenStudy (anonymous):

1. \[c=10+\frac{ 5 }{ 2 }d\] 2. \[d=\frac{ 2 }{ 5 }x-4\] 3. \[p=-\frac{ 2 }{ 3 }q-3\] 4. \[q=-\frac{ 3 }{ 2 }p-\frac{ 9 }{ 2 }\] 5. \[x=3+\frac{ 1 }{ 2 }y\] 6. \[y=2x-6\] 7. \[b=-\frac{ 4 }{3 }a+\frac{ 8 }{ 3 }\] 8. \[a=\frac{ 3 }{ 2 }bc-\frac{ 9 }{ 2 }\] 9. \[b=\frac{ -9-2a }{ -3c }\] 10. \[c=\frac{ -9-2a }{ -3b }\]

jabez177 (jabez177):

THANK YOU SO MUCH!!!!! Here's your medal!

OpenStudy (anonymous):

No problem! :)

jabez177 (jabez177):

Can you help me with some slope stuff? I have only 6 questions of those. I'll tag you again and give you some more medals. :)

OpenStudy (anonymous):

Yeah! I have to leave soon but I'll answer as many as I can